OpenStudy (rebeccapink6398):

medal+fan if you help meh Sharon is the youngest in her family. Her sister Anne is 3 years older than she is. Her brother Tom is 5 years older than Anne. The sum of their ages is 62. How old is Sharon?

OpenStudy (anonymous):

\((s)+(s+3)+(s+3+5)=62\)

OpenStudy (anonymous):

let x = Sharon's age Anne = x + 3 Tom = x + 5 equation: x + x + 3 + x + 5 = 62

OpenStudy (rebeccapink6398):

i was thinking mayeb she is 8

OpenStudy (anonymous):

that get u \(3s+11=62\)

OpenStudy (anonymous):

Can you solve the equation from there?

OpenStudy (anonymous):

solve that u can get the answer

OpenStudy (rebeccapink6398):

i dont get what you said molly

OpenStudy (anonymous):

if u put what i said and what he said together it'll make more sense

OpenStudy (anonymous):

i just simplified his equation that's all

OpenStudy (rebeccapink6398):

i just don't understand how to even get the answers from your guys equation

OpenStudy (anonymous):

oh u see... if you mix around the numbers... x+x+x+3+3+5 right?

OpenStudy (anonymous):

x+x+x=3x right? 3+3+5=11 yes?

OpenStudy (anonymous):

so it gives you 3x+11 light?

OpenStudy (rebeccapink6398):

what gives you 3x? like where did you get the second 3 from!?

OpenStudy (anonymous):

oh.... 3x means there are three x'es

OpenStudy (anonymous):

that's where the 3 came from

OpenStudy (rebeccapink6398):

so 3 times the peoples ages then plus 11? and then take away 62?

OpenStudy (anonymous):

also written as \(3\times x\)

OpenStudy (anonymous):

no. 3x+11 is supposed to be equal to 62.

OpenStudy (anonymous):

62 is the sum of their ages

OpenStudy (rebeccapink6398):

so its 3*saras age,toms age,annes age? and then plus 11. where di you get the 11 from?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ok u see.... let's say sara's age is s. anne's age is 3 more than sara's so it's s+3

OpenStudy (anonymous):

do you get this part?

OpenStudy (anonymous):

and tom is 5 years older than anne, who is 3 years older than sara, so that gives you s+3+5. yes?

OpenStudy (rebeccapink6398):

yes

OpenStudy (anonymous):

so all of them adds up to 62. so that gives you s+s+3+s+3+5=62. do you get that part?

OpenStudy (anonymous):

(s)+(s+3)+(s+3+5) sara+anne+ tom

OpenStudy (anonymous):

yes?

OpenStudy (rebeccapink6398):

ok

OpenStudy (anonymous):

and so that gives you 3s+11 as i said...

OpenStudy (anonymous):

do you get that part?

OpenStudy (rebeccapink6398):

yes :)

OpenStudy (anonymous):

3s+11=62

OpenStudy (anonymous):

no now it's just an equation for you to solve.

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

do you know how to solve that eq

OpenStudy (anonymous):

finding the x?

OpenStudy (rebeccapink6398):

no i still dont get it DX

OpenStudy (rebeccapink6398):

wait when you say 3s it means 3 times s. you know that right? :)

OpenStudy (rebeccapink6398):

i was tought that